A 0.320-kg block of a pure material is heated from 20.0°C to 65.0°C by the addition of 2.88 kJ of energy. Calculate its specific heat. I got 200 but that is not the right answer and it needs to be in kJ/(kg · °C)
Answer #1

Heat added,

Q = m*c*delta_T

delta_T=65-20 =45 C

m=0.320 kg,

Q=2.88 *10^3 J

2880=0.320*c*45

c=200 J/kg*k

=0.2 KJ/kg*k (Specific heat is 0.2Kj/kgk not 200 KJ/kgk)